Crew 216 Operations Report 07Dec2019

[title Operations Report – December 7th]

[category operations-report]

Crew 216 Operations Report 07-Dec-19
SOL: 13
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 12
From what time last night: 1905
To what time this morning: 0700
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 73%
Diesel Reading – 31
Station Propane Reading – 49
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 14656315
Water (static tank) (gal): 450 (filled today, 2 trips to Hanksville)
Static to Loft Pump used – yes
Water in Green Hab (gal): 238
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no. Rover cleaned for crew turn-over.
Hours: 116.4
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging: yes
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no Rover cleaned for crew turn-over.
Hours: 121.4
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where? Yes, water re-supply, 2 trips to Hanksville.
CrewCar used and why, where?
General notes and comments: Simulation ended 062100 Dec 2019. Station cleaning continues. Food supply re-stocked.
Water supply re-stocked.
Summary of internet: Erratic/unstable wifi connection.
Summary of suits and radios: Suit 4 airflow is weak, possibly a battery charging issue.
Summary of Hab operations: HAB cleaning continues. Hot water line vibration is extreme when pump
engages, recommend securing it to HAB wall or risk fitting failure.
Summary of GreenHab operations: Routine watering. Greenhab cleaned for crew turn-over.
Summary of Science Dome operations: Lab cleaned for crew turn-over.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved. RAM cleaned for crew turn-over.
Summary of any observatory issues: Observatory cleaned for crew turn-over.
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: NA

Crew 216 Astronomy Report 07Dec2019

[title Astronomy Report – December 7th]

[category astronomy-report]

Crew 216 Astronomer Report 07-DEC-19
Crew Astronomer: Michael Ho

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: none
Objects to be Imaged this Evening: none
Images submitted with this report: none
Problems Encountered: nil

MUSK OBSERVATORY
Solar Features Observed: None.
Images submitted with this report: nil
Problems Encountered: None.

END OF MISSION PHOTOGRAPHS FOR MUSK OBSERVATORY:
4 attached. (Apologies the one with the complete power cable was just taken in the dark Musk Observatory).

Crew 216 Science Report 07Dec2019

[title Science Report – December 7th]

[category science-report]

Crew 216 Research Report 7-Dec-2019 Final
Sol 13
Submitted by: Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Final surveys completed. Results to be analyzed post-mission.

2) Medical Casts – Nothing further to report. Data to be analyzed post-mission.

3) In-Situ Fluorescent Mineral Prospecting – Nothing further to report.

4) MDRS Mapping – Crew Commander discussed mapping with director. Actions to be followed up post-mission.

5) MDRS Communications – Nothing further to report. Follow up report will be developed if desired.

Crew 216 GreenHab Report 07Dec2019

[title GreenHab Report – December 7th]

[category greenhab-report]

Crew 216 GreenHab Report 07-DEC-19

Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 11.50 am, closed at 5.00 pm(?) – thanks to the Martian elves.
Average temperature: 20C
Hours of supplemental light: Automatic system
Daily water usage for crops: 12L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 9.30 am; 6.45 pm.

SEASONAL PLANTING HISTORY:
(Plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)
(Plants introduced since 23/11/2019 till 07/11/2019 please refer to Crew 216 GreenHab Report of 06/12/2019.)

OTHER NOTES:
Harvest: None
Support/supplies requested:nil.

Crew 216 Sol Summary 07Dec2019

[title Sol Summary – December 7th]

[category sol-summary]

Crew 216 Sol Summary Report 07-DEC-19
Sol: 13
Summary Title: Departure prep
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: Cleaning and packing for departure in the morning.
Look Ahead Plan: Saying good-bye to MDRS.
Anomalies in work: N/A.
Weather: Cloudy and cold overnight.
Crew Physical Status: Healthy
EVAs: None
Reports to be filed: Sol Summary Report, Operations Report, GreenHab Report, Research Report, and Astronomy Report.
Support Requested: None